OpenStudy (anonymous):

I need help with some algebra homework. It is simplifying expressions with exponents. Please help!!

OpenStudy (anonymous):

\[-2n ^{4} (12n ^{-7})\]

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

Multiply the coefficients and add the exponents \[-2n^4(12n^-7)=(-2 \times 12)n^{4+(-7)}\]

OpenStudy (anonymous):

but i need to simplify it. example: \[-3n ^{2} (8n ^{-5}) =\frac{ -24 }{ n ^{3} } \]

OpenStudy (anonymous):

I'm having trouble understanding the whole steps...

OpenStudy (anonymous):

-24/n³ is simplifed

OpenStudy (anonymous):

To multiply expressions when the bases are the same, add the exponents and multiply the coefficients \[(-3n^2)(8n^{-5})=(-3 \times 8)n^{2+(-5)}=-24n^{-3}\] For negative exponents, if you take the reciprocal of the base, you can make the exponent positive Rule: \[x^{-m}=\left( \frac{ 1 }{ x } \right)^m\] That's why the -3 became 3 and the n moved to the denominator \[-24n^{-3}=-24\left( \frac{ 1 }{ n } \right)^3=\frac{ -24 }{ n^3 }\]

OpenStudy (anonymous):

That was exactly what my teacher was trying to show us. It's just I was having trouble. Thanks for the help! :)

OpenStudy (anonymous):

you're welcome